OpenStudy (anonymous):

i need some assistance with this one. I have a chart with x & y parts of an ordered pair. I got 2 of them having trouble with the last one. 6x-2y=12 6x-2(-12)=12 right?

OpenStudy (anonymous):

6x-2y=12 dividing the equation with 2 we get 3x-y=6 now taking x=0 we get y=-6 so first ordered pair is (0,-6) taking y as -12 we get 3x=6-12=-6 or x=-2

OpenStudy (anonymous):

i am so sorry I forgot to tell you that y=-12